OpenStudy (anonymous):

Could someone help me on this question please, Simplify: [2root5 + 3root2] / [5root5 + 2root2]

OpenStudy (anonymous):

38/117

OpenStudy (anonymous):

The answer is 38+11root10 / 117 but im not sure how to get that :S

OpenStudy (anonymous):

just guesss

OpenStudy (anonymous):

lol i rather not :P

OpenStudy (anonymous):

I am assuming you intend:\[\dfrac{2\sqrt 5 + 3\sqrt 2}{5 \sqrt 5 + 2 \sqrt 2}\]Multiply by the conjugate of the denominator over the conjugate of the denominator:\[\dfrac{2\sqrt 5 + 3\sqrt 2}{5 \sqrt 5 + 2 \sqrt 2}\cdot \dfrac{5 \sqrt 5 - 2 \sqrt 2}{5 \sqrt 5 - 2 \sqrt 2}\]\[\dfrac{10(5)-4\sqrt{10} + 15 \sqrt {10}-6(2)}{25(5)-4(2)}\]Thus, the answer is\[\dfrac{38+11\sqrt{10}}{117}\]

OpenStudy (anonymous):

yeay yakeglee answer best

OpenStudy (anonymous):

Ohh I get it thanks so much for your help

OpenStudy (anonymous):

Give me medals then :P

OpenStudy (anonymous):

medal and i'll become a fan :D